### v1

# Technical insights and legal considerations for advancing federated learning in bioinformatics

Abstract
Federated learning leverages data across institutions to improve clinical discovery while complying with data-sharing restrictions and protecting patient privacy. As the evolution of biobanks in genetics and systems biology has proved, accessing more extensive and varied data pools leads to a faster and more robust exploration and translation of results. More widespread use of federated learning may have the same impact in bioinformatics, allowing access to many combinations of genotypic, phenotypic and environmental information that are undercovered or not included in existing biobanks. This paper reviews the methodological, infrastructural and legal issues that academic and clinical institutions must address before implementing it. Finally, we provide recommendations for the reliable use of federated learning and its effective translation into clinical practice.
keywords:
Federated Machine Learning, Exposome, Secure distributed analysis, Data privacy, Collaborative genomics1 Introduction
Sharing personal information has been increasingly regulated in both the EU (with the GDPR and the AI act; European Union, 2016, 2024) and the US (with the National AI Initiative Act; U.S. Congress, 2020) to mitigate the personal and societal risks associated with their use, particularly in connection with machine learning and AI models (Cath et al., 2018). These regulations make multi-centre studies and similar endeavours more challenging, impacting biomedical and clinical research.
Federated learning (FL; McMahan et al., 2017; Ludwig and Baracaldo, 2022) is a technical solution intended to reduce the impact of these restrictions. FL allows multiple parties to train a global machine learning model collaboratively from the respective data without sharing the data themselves and without any meaningful model performance degradation. Instead, parties only share model updates, making it impractical to reconstruct personal information when the appropriate secure computational measures are implemented (Wainakh et al., 2022).
This approach strengthens security by keeping sensitive information local, improves privacy by minimising data exposure even between the parties involved, and limits risk of data misuse by allowing each party to retain complete control over its data (Truong et al., 2021). If enough parties are involved, FL may potentially access larger and more varied data pools than centralised biobanks can provide, particularly if there are legal (or other) barriers to data centralisation, resulting in more accurate and robust models than those produced by any individual party.
FL has proven to be a valuable tool for biomedical research expected to gain further traction in the coming years. Its use has improved breast density classification models (accuracy up by 6%, generalisability up by 46%; Roth et al., 2020), COVID-19 outcome prediction at both 24h and 72h (up 16% and 38%; Dayan et al., 2021) and rare tumour segmentation (up by 23-33% and 15%; Pati et al., 2022) compared to single-party analyses. A consortium of ten pharmaceutical companies found that FL improved structure-activity relationship (QSAR) models for drug discovery (both up 12% Heyndrickx et al., 2023). Early-stage applications building predictive models from electronic health records (Brisimi et al., 2018) have also confirmed no practical performance degradation compared to pooling data from all parties.
To achieve such results, a real-world implementation of FL must overcome several methodological, infrastructural and legal issues. However, biomedical FL literature reviews (Xu et al., 2020; Chowdhury et al., 2022, among others) are predominantly high-level and considered simulated rather than real-world implementations. Here, we will cover federated methods designed explicitly for bioinformatics and discuss the infrastructure they need and how they satisfy the legal requirements.
To this end, we have structured the remainder of the paper as follows: We first review the fundamental concepts and design decisions of FL in Section 2, including different topologies (Section 2.1), hardware and software (Section 2.2), data layouts in different parties (Sections 2.3 and 2.4), security (Section 2.5) and privacy concerns (Section 2.6). In Section 3, we contrast and compare bioinformatics FL methods for proteomics and differential expression (Section 3.1), genome-wide association studies (GWAS; Section 3.2), single-cell RNA sequencing (Section 3.3), multiomics (Section 3.4) and medical imaging (Section 3.5) applications. We conclude the section with notable examples of ready-to-use software tools (Section 3.6). Section 4 provides examples of federated operations common in bioinformatics. Finally, we discuss the relevant legal implications of using FL (Section 5) before discussing our perspective on its reliable use and translation to clinical practice (Section 6).
2 Federated learning
FL is a collaborative approach to machine learning model training, where multiple institutions form a consortium to jointly train a shared model by exchanging model updates rather than raw patient data. Typically, FL involves data holders (called “clients”) sharing their local contributions with a server (McMahan et al., 2017) as outlined in Figure 1. The server then creates and shares back a global model, inviting the data holders to update and resubmit their contributions. This process is iterative and involves several rounds of model update exchanges. Unlike traditional centralised computing, FL does not store patient data in a central location. Instead, patient data remain under the control of the respective data owners at their sites, enhancing privacy.
FL has similarities with distributed computing, meta-analysis, and trusted research environments (TREs) but also has key differences, which we highlight below.
Distributed computing (DC) (Zomaya, 2006) divides a computational task among multiple machines to enhance processing speed and efficiency. Usually, DC starts from a centrally managed data set spread across different machines, which is assumed to contain independent and identically distributed observations. Each machine is tasked to process a comparable quantity of data. In contrast, clients independently join FL with their locally-held data, which may vary significantly in quantity and distribution. While sharing some techniques with FL, distributed computing aims for computational efficiency and lacks its privacy focus.
On the other hand, meta-analyses (Toro-Domínguez et al., 2021) aggregate results across previously completed studies using statistical methods to account for their variations, thus allowing researchers to synthesise findings without accessing raw data and preserve the privacy of individual data sets. Here, FL collaboratively trains a joint model using distributed data to iteratively update it while meta-analysis constructs it in a single step from the pre-existing results. Multiple studies on sequencing data have demonstrated that FL produces results closer to centralised analysis than from meta-analysis (Mendelsohn et al., 2023; Zolotareva et al., 2021).
TREs (Kavianpour et al., 2022) provide access to data within a controlled, secure computing environment for conducting analyses, almost always disallowing data sharing. Some TREs have a centralised data location and governance; an example is the Research Analysis Platform (RAP), the TRE for the UK Biobank (UKB; Sudlow et al., 2015). Others, such as FEGA (Federated European Genome-phenome Archive, 2024), are decentralised. Each institution maintains its data locally; only the relevant data are securely transferred to the computing environment when the analysis is authorised. Unlike FL, the learning process is not distributed across the data holders. Thus, the tradeoff between TREs and FL is between a centralised, trusted entity with large computational facilities that can place substantial restrictions on the analysis, and a consortium that requires all parties to apply governance guidelines and provide compute, but can scale both data access and privacy guarantees.
2.1 Topologies
The topology of the FL consortium is determined by the number of participating parties and their defined interactions. Some examples are illustrated in Figure 2. The most common is the centralised topology, where multiple data-holding parties (the clients) collaboratively train a shared machine learning model through a central server (the aggregator) that iteratively collects model updates from each client, updates the global model, and redistributes it back to the clients. Typically, clients do not communicate directly; they only communicate with the central server. In contrast, a decentralised topology (Beltrán et al., 2023) lacks a dedicated aggregation server. All consortium parties can potentially be model trainers and aggregators, interacting via peer-to-peer communication. Hybrid configurations include, for instance, using two servers: one server handles aggregation of noisy local models, while the other performs auxiliary tasks, such as noise aggregation (Nasirigerdeh et al., 2021). Clients can communicate with the servers, and servers can communicate between themselves, but clients cannot communicate with each other.
We will focus on the standard centralised topology and its two-server variant here because, to our knowledge, no bioinformatics applications use decentralised topologies.
2.2 Hardware and software
Hardware, software and models should be chosen with knowledge of the data and inputs from domain and machine learning specialists to design an effective machine learning pipeline (Scutari and Malvestio, 2023).
In terms of infrastructure, FL requires computational resources for each client and server. The optimal hardware configuration depends on the models to be trained; at a minimum, each client must be able to produce model updates from local data, and each server must be able to aggregate those updates and manage the consortium. Connection bandwidth is not necessarily critical: to date, client-server communications contain only a few megabytes of data, reaching 150MB only for large computer vision models, and can be made more compact through compression and model quantisation (Camajori Tedeschini et al., 2022). On the other hand, latency may be a bottleneck if it limits the hardware utilisation.
As for software, several dedicated FL frameworks, many of which are comparatively analysed in (Riedel et al., 2024), provide structured tools and environments for developing, deploying, and managing federated machine learning models. While some frameworks, such as Tensorflow Federated (TFF; Google, 2024), specialise in particular models, others support a broader range of approaches. Notable open-source examples include PySyft (Ziller et al., 2021) and Flower (Beutel et al., 2020). Both are supported by active communities and integrate with PyTorch to train complex models. PySyft is a multi-language library focusing on advanced privacy-preserving techniques, including differential privacy and homomorphic encryption. Flower is an FL framework: its modular design and ease of customisation make it particularly useful for large-scale and multi-omics studies involving heterogeneous devices and clients. We will provide examples using these frameworks in Section 3 before discussing frameworks explicitly designed for bioinformatics in Section 3.6.
Other frameworks target healthcare and biomedical applications but not bioinformatics specifically. For instance, OpenFL (Foley et al., 2022) is designed to facilitate FL on sensitive EHRs and medical imaging data; it supports different data partitioning schemes (Section 2.4) but struggles with heterogeneous cross-device FL (Section 2.3). NVIDIA Clara, which was used in Dayan et al. (2021), has similar limitations.
2.3 Usage scenarios: cross-device and cross-silo
FL applications take different forms in different domains. Many small, low-powered clients, such as wearable medical devices from the Internet of Things, may produce the data needed to train the federated machine learning model. Such cross-device communications are often unreliable: passing lightweight model updates instead of raw data largely addresses connectivity issues and privacy risks.
FL may also involve a small number of parties, each possessing large amounts of sensitive data (Huang et al., 2022), stored within their “data silos”. In this cross-silo scenario, common in healthcare and bioinformatics, the priority is to minimise the privacy risks associated with data sharing and comply with regulations. Minimising large data transfers is also computationally advantageous when modelling large volumes of information, such as whole-genome sequences.
These two scenarios differ in how they handle model updates. In the cross-silo scenario, all (few) data holders in the consortium must participate in each update. In contrast, we can rely on a subset of (the many) data holders in the cross-device scenario because each holds a smaller share of the overall data. This article focuses on the cross-silo scenario, as nearly all bioinformatics applications fall within this framework.
2.4 Data partitioning and heterogeneity
Data may be partitioned along two axes: each party may record the same features for different samples or features describing the same samples (Figure 3). In the first scenario, known as horizontal FL,111This naming convention assumes that samples (features) are the rows (columns) of a tabular data set. different parties may each possess genomic sequencing data from different individuals. In contrast, in vertical FL, one party may hold data from one omic type (say, genomic data), while another may have data from a different phenotype or omic type (say, proteomic data) for the same individuals. Horizontal FL is by far the most prevalent approach in bioinformatics.
Significant variations in sample size and feature distributions between data holders often exist. This heterogeneity allows FL to better capture the variability of the underlying population, resulting in transferrable models that generalise well (Sheller et al., 2020). Clearly, if data holders collect observations from distinct populations, any federated model trained from them must be correctly specified to capture population structure and avoid bias in inference and prediction. If the populations are known, we can train targeted population-specific models alongside the global one (Tan et al., 2022). Otherwise, we can use clustering to identify them from the available data (Sattler et al., 2021). Accounting for variations in measurements, definitions and distributions to harmonise data across parties is also fundamental but is much more challenging because access to data is restricted, even more so than in meta-analysis (Camajori Tedeschini et al., 2022).
2.5 Security and privacy
FL reduces some privacy and security risks by design by passing model updates between parties instead of centralising data in a single location. However, it does not eliminate them completely.
In terms of privacy, deep learning models are the most problematic in machine learning because of their ability to memorise training data. They leak individual observations during training (through model updates; Geiping et al., 2020), after training (through their parameters; Haim et al., 2022) and during inference (membership attacks; Shokri et al., 2017; Hu et al., 2022). More broadly, individual reidentification is an issue for genetic data (Homer et al., 2008) and all the models learned from them. For instance, (Cai et al., 2015) has demonstrated that it is possible to identify an individual from the linear model learned in an association study from just 25 genes. However, such works make unrealistic assumptions on the level of access to the models and the data (Kolobkov et al., 2024): even basic infrastructure security measures and the distributed nature of the data will make such identification difficult under the best circumstances. The privacy-enhancing techniques discussed in Section 2.6 can make such efforts completely impractical.
As for security, we must consider different threat models, understanding what information requires protection, their vulnerabilities, and how to mitigate or respond to threats. Internal and external threats to the consortium should be treated equally with security in depth design and implementation decisions that consider parties untrusted. Security threats, such as membership attacks and model inversion attacks (Fredrikson et al., 2015), can originate equally from parties and external adversaries that seek to abuse the model inference capabilities to extract information about the data. On the other hand, adversarial attacks are more likely to originate from consortium parties that seek to introduce carefully crafted data or model updates into the training process to produce a global model with undesirable behaviour. Some examples are data poisoning (Sun et al., 2022b), manipulation (Jagielski et al., 2018) and Byzantine attacks (Li et al., 2023a).
Encrypting communication channels, implementing strict authentication (to verify each party’s identity) and authorisation (to control which information and resources each party has access to or shares) schemes, and keeping comprehensive access logs for audit can secure any machine learning pipeline, including federated ones. Similarly, using an experiment tracking platform makes it possible to track data provenance, audit both the data and the training process and ensure the reproducibility of results (Scutari and Malvestio, 2023). These measures must be complemented by federated models resistant to these threats at training and inference time, as thoroughly discussed in Yin et al. (2021).
2.6 Privacy-enhancing techniques
Privacy-enhancing techniques improve the confidentiality of sensitive information during training. We summarise the most relevant below, illustrating themin Figure 4.
Homomorphic encryption (HE; Gentry, 2009) is a cryptographic technique that enables computations to be performed directly on encrypted data (ciphertexts) without requiring decryption. The outcome of operations on ciphertexts matches the result of performing the same operations on the corresponding non-encrypted values (plaintexts) when decrypted. HE can be either fully homomorphic (FHE), which allows for arbitrary computations, or partially homomorphic (PHE), which supports only a specific subset of mathematical operations. For instance, the Paillier PHE scheme (Paillier, 1999) only supports additive operations on encrypted data. FHE requires considerable computational resources for encryption and decryption. PHE is less flexible but computationally more efficient, making it a common choice in practical applications.
Secure multiparty computation (SMPC; Zhao et al., 2019) is a peer-to-peer protocol allowing multiple parties to compute a function over their data collaboratively, similarly to Figure 2 (centre). Each data holder divides their data into random shares and distributes them among all parties in the consortium, thus ensuring that no single party can access the complete data set. The shares are then combined during the computation process, often with the assistance of a server, to produce the correct result while preserving data privacy. SMPC ensures high security with exact results and keeps data private throughout the computation process. However, SMPC is computationally intensive and requires peer-to-peer communication, leading to high communication overhead. Its complexity also increases with the number of participants, limiting scalability.
Another approach to securing FL is using an aggregator and a compensator server in a centralised two-server topology (Figure 2, right; Nasirigerdeh et al., 2021). Each client adds a noise pattern to their local data, sharing the former with the compensator (which aggregates all noise patterns) and the latter with the aggregator (which aggregates the noisy data and trains the model). The aggregator then obtains the overall noise pattern from the compensator and removes it from the aggregated noisy data, allowing for denoised model training. This two-server approach is efficient: it requires neither extensive computation in the clients nor peer-to-peer communication. However, it makes infrastructure more complex and requires trust in both servers not to collude to compromise the privacy of individual contributions.
Unlike the above methods, which are encryption-based methods ensuring data confidentiality during transmission or storage, differential privacy, another popular technique for data-protection in federated learning, is not an encryption system but rather a technique that focuses on privacy by ensuring that the output of data analysis does not leak sensitive information about the underlying dataset. Differential privacy (DP; Ficek et al., 2021) achieves this by a mathematical framework designed to enable analyses to remain statistically consistent regardless of whether any specific individual’s data is included or excluded. This property guarantees that sensitive information about individuals cannot be inferred from the results up to a preset “privacy budget” worth of operations. DP is typically implemented by introducing noise into the data (Schein et al., 2019; Cai et al., 2021), weight clipping in the training process (Abadi et al., 2016; Jayaraman and Evans, 2019) or predictions (Nissim et al., 2007; Dwork and Feldman, 2018) to obfuscate individual contributions. The amount of noise must be carefully calibrated to balance predictive accuracy and privacy within the analysis: too little noise undermines privacy, and too much reduces performance. This effect is more pronounced within specific subgroups underrepresented in the training set (Bagdasaryan et al., 2019).
3 Federated learning in bioinformatics
Most FL literature focuses on general algorithms and is motivated by applications other than bioinformatics, such as digital twins for smart cities (Ramu et al., 2022), smart industry (Zhang et al., 2021) and open banking and finance (Long et al., 2020). Even the clinical literature mainly focuses on different types of data and issues (Dayan et al., 2021; van Rooden et al., 2024). Here, we highlight and discuss notable examples of FL designed specifically for bioinformatics. They are all in the early stages of development, so their reliability, reproducibility, and scalability are open questions. However, they hint at the potential of FL to perform better than meta-analysis and single-client analyses on real-world data, comparing favourably to centralised data analyses in which data are pooled in a central location while addressing data sharing and use concerns (Zolotareva et al., 2021; Xu et al., 2020).
3.1 Proteomics and differential gene expression
Proteomics studies the complex protein dynamics that govern cellular processes and their interplay with physiological and pathological states, such as cancer (Maes et al., 2015), to improve risk assessment, treatment selection and patient monitoring. Differential expression analyses focus specifically on comparing expression levels across different conditions, tissues, or cell types to identify genes with statistically significant differences (Rodriguez-Esteban and Jiang, 2017).
In addition to the issues discussed in Section 2, FL in proteomics must overcome the challenge of integrating data from different platforms (Rieke et al., 2020) while accounting for imbalanced samples and batch effects. Cai et al. (2022) produced a federated implementation of DEqMS (FedProt; Zhu et al., 2020) for variance estimation in mass spectrometry-based data that successfully identifies top differentially-abundant proteins in two real-world data sets using label-free quantification and tandem mass tags.
Zolotareva et al. (2021) implemented a federated limma voom pipeline (Law et al., 2014) on top of HyFed (Nasirigerdeh et al., 2021), which uses the aggregator-compensator two-server topology we described earlier. This approach was showcased on two extensive RNA-seq data sets, proving robust to heterogeneity across clients and batch effects. Hannemann et al. (2024) trained a federated deep-learning model for cell type classification using both Flower and TFF and different architectures, with similar results.
3.2 Genome-wide association studies
Genome-wide association studies (GWAS) aim to identify genomic variants statistically associated with a qualitative (say, a case-control label) or quantitative trait (say, body mass index). These studies mainly use regression models, which can be largely trained using general-purpose federated regression implementations with minor modifications to address scalability and correct for population structure (see, for instance, Kolobkov et al., 2024).
Li et al. (2022) has developed the most complete adaptation of these models to federated GWAS in the literature: it provides linear and logistic regressions with fixed and random effects and accounts for population structure via a genomic relatedness matrix. Wang et al. (2022) further provides a federated estimator for the genomic relatedness matrix. Finally, Li et al. (2024) describes the federated association tests for the genomic variants associated with this model. All these steps incorporate HE to ensure privacy in the GWAS.
As an alternative, Cho et al. (2024) build on REGENIE (Mbatchou et al., 2021) to avoid using a genomic relatedness matrix and increase the scalability of GWAS while using MPC and HE to secure the data. Despite the overhead introduced by the encryption, this approach is efficient enough to work on a cohort of 401k individuals from the UK Biobank and 90 million single-nucleotide polymorphisms (SNPs) in less than 5 hours.
3.3 Single-cell RNA sequencing
Single-cell RNA sequencing (scRNA-seq) measures gene expression at the cellular level rather than aggregating it at the tissue level as in bulk RNA sequencing, identifying the distinct expression profiles of individual cell populations within tissues (Hwang et al., 2018; Papalexi and Satija, 2018).
Wang et al. (2024) developed scFed, a unified FL framework integrating four algorithms for cell type classification from scRNA-seq data: the ACTINN neural network (Ma and Pellegrini, 2020), explicitly designed for this task; a linear support vector machine; XGBoost based on Li et al. (2023b); and the GeneFormer transformer (Theodoris et al., 2023). They evaluated scFed on eight data sets evenly distributed among 2–20 clients, suggesting that the federated approach has a predictive accuracy comparable to that obtained by pooling the data and better than that in individual clients. However, the overhead during training increases with the number of clients, limiting the scalability to larger consortia.
3.4 Multi-omics
Proteomics, genomics, and transcriptomics capture different aspects of biological processes. Integrating large data sets from different omics offers deeper insights into their underlying mechanisms (Civelek and Lusis, 2014). Vertical FL allows multiple parties to combine various features of the same patients into multimodal omics data sets without exposing sensitive information (Liu et al., 2024). For instance, Wang et al. (2023) trained a deep neural network with an adaptive optimisation module for cancer prognosis evaluation from multi-omics data. The neural network performs feature selection while the adaptive optimisation module prevents overfitting, a common issue in small high-dimensional samples (Rajput et al., 2023). This method performs better than a single-omic analysis, but the improvement in predictive accuracy is strongly model-dependent. Another example is Danek et al. (2024), who built a diagnostic model for Parkinson’s disease: they provided a reproducible setup for evaluating several multi-omics models trained on pre-processed, harmonised and artificially horizontally federated data using Flower. Their study identifies a general but not substantial reduction in FL performance compared to centrally trained models, which increases with the number of clients and is variably affected by client heterogeneity.
3.5 Medical imaging
Medical imaging studies the human body’s interior to diagnose abnormalities in its anatomy and physiology from digital images such as those obtained by radiography, magnetic resonance and ultrasound devices (Suetens, 2017). It is the most common application of FL in the medical literature (Chowdhury et al., 2022). As a result, protocols for image segmentation and diagnostic prediction are well documented. Notable case studies target breast cancer (Roth et al., 2020), melanomas (Haggenmüller et al., 2024), cardiovascular disease (Linardos et al., 2022), COVID-19 (Yang et al., 2021; Dayan et al., 2021).
Machine learning applications that use medical imaging data typically face challenges such as incomplete or inaccurate labelling and the normalisation of images from different scanners and different protocols. Bdair et al. (2022) explored a federated labelling scheme in which clients produced ground-truth labels for skin lesions in a privacy-preserving manner, improving classification accuracy. Yan et al. (2023) also proposed an efficient scheme to use data sets mainly comprising unlabelled images, focusing on chest X-rays. Furthermore, Jiang et al. (2022) apply FL to learn a harmonised feature set from heterogeneous medical images, improving both the classification and segmentation of histology and MRI scans.
3.6 Ready-to-use FL tools for bioinformatics
The need for user-friendly FL implementations of common bioinformatics workflows has driven the creation of secure collaborative analysis tools (Berger and Cho, 2019; Froelicher et al., 2021; Wan et al., 2022). Two notable examples are sfkit and FeatureCloud.
The sfkit framework (Mendelsohn et al., 2023) facilitates federated genomic analyses by implementing GWAS, principal component analysis (PCA), genetic relatedness and a modular architecture to complement them as needed. It provides a web interface featuring a project bulletin board, chat functions, study parameter configurations and results sharing. State-of-the-art cryptographic tools for privacy preservation based on SMPC and HE ensure data protection (Mouchet et al., 2020).
FeatureCloud (Matschinske et al., 2023) is an integrated solution from which end users without programming experience can build custom workflows. It provides modules to run on the clients and servers in the consortium. Unlike sfkit, FeatureCloud allows users publish applications in its app store, including regression models, random forests and neural networks. Developers must also document how privacy guarantees are implemented in their apps.
4 How to conduct federation of specific operations in bioinformatics
This section provides practical insights to help a reader interested in building a federated and secure analogue of an existing bioinformatics algorithm. We focus on horizontal FL with the centralised topology from Figure 2 (left). Consider different clients, each possessing a local data set , where . Each data set contains samples, denoted as , where represents the sample index, and represents the features for each sample. We denote a row (column) of the matrix as (). This describes a distributed data set of observations:
The following sections assume that an FL consortium has been established, the necessary infrastructure is operational, and an appropriate FL framework has been selected and installed. It is also assumed that a secure aggregation protocol has been chosen, such as those described in Section 2.6 and Figure 4. The choice of a specific secure aggregation protocol may depend on several factors, including technology and infrastructure (e.g., the availability of a specific FL topology that drives the choice), privacy risks, or scalability concerns, as discussed in Section 2.6. In the following sections, we provide a general overview of sum-based mathematical operations built upon a secure aggregation protocol, as well as operations involving federated averaging (FedAvg; McMahan et al., 2017).
Coding examples using Flower (Beutel et al., 2020) are available in our GitHub repository (https://github.com/IDSIA/FL-Bioinformatics). We chose Flower because it has a shallow learning curve for new FL users and provides a good balance between simplicity and flexibility when implementing custom algorithms. Riedel et al. (2024) also identified Flower as a promising framework because it has a large, active, and growing community of developers and scientists and because of the extensive tutorials and documentation. In our examples, secure summation is performed using the secure aggregation protocol SecAgg+ (Bell et al., 2020). This protocol combines encryption with SMPC, using a multiparty approach in which each client interacts with only a subset of the others. It is particularly suitable for several FL contexts, as it is robust to client dropout and highly scalable. In particular, a relevant aspect of the bioinformatics domain is that it scales linearly with the size of the vectors to be aggregated (Li et al., 2021).
4.1 Sum-based computations
Let be real numbers stored by individual clients. We define the secure sum of these numbers, performed through the selected secure aggregation protocol, as . We can build on this simple secure sum to construct a wide range of operations. However, note that as the complexity of operations increases, the amount of information revealed to the server may also increase. Sum-based operations include:
The overall sample size of the distributed data set as from the local sample sizes .
The mean value of the -th feature, given , as
Each client computes the inner sum on their local data, whereas the outer one is a secure sum aggregated across clients by the server.
The variance of the -th feature, given and , as
which can be used to standardise the -th feature as .
The Pearson correlation coefficient of two features and , given and , as
The matrix , as , where is a secure element-wise sum. This matrix is equivalent to the covariance matrix for standardised data sets and is commonly used for PCA.
Beyond these general-purpose examples, many operations specific to bioinformatics pipelines also rely on simple sums. These operations are often straightforward generalisations or compositions of the examples introduced above.
In differential gene expression studies, for instance, filtering out weakly expressed genes is standard practice. Weakly expressed genes can be defined as those whose expression values fall below a specified threshold in, for instance, 70% of the samples. Let be a vector belonging to client , where each vector component represents the number of samples in which the expression level of the gene (e.g., counts) exceeds the threshold . The server can securely calculate and identify weakly expressed genes as those whose corresponding components of are smaller than 0.7.
A fundamental preliminary step in a GWAS is identifying the minor allele and its frequency. Let , , , and be vectors belonging to client , where each component corresponds to a specific SNP. The components of , , , represent the number of samples in which nucleotides , , , are observed, respectively. The server can securely compute the aggregated allele counts across all clients as and similarly , , (where can also be computed by difference from and the other three vectors). For each SNP, the minor allele is determined by comparing the corresponding components of , , , : the allele with the smaller value is designated as the minor allele. This operation is crucial because the minor allele within a single client’s population may differ from the minor allele when considering the whole distributed data set. Ensuring a consistent definition of the minor allele across all clients is essential for reliable downstream analyses.
4.2 Federated averaging computations
FedAvg is a widely used algorithm for training deep neural networks in FL. It iteratively computes a weighted average of model parameters across clients, with weights proportional to the local sample sizes . Thus, it can be applied to any parametric model, including linear models.
FedAvg proceeds as illustrated in Figure 1. The server first broadcasts an initial global model with parameters . At each step of the algorithm, clients start with the global model and perform local updates to produce updated local models . The global model is updated after each round of local training as the weighted sum of the local models:
where we use the secure sum for aggregation (FedAvg is itself a sum-based operation). After aggregation, the updated global model is distributed back to the clients.
However, many bioinformatics pipelines rely on linear models rather than deep learning models. One commonly used model is logistic regression, which is applied in tasks such as gene expression analysis and GWAS. A federated implementation of logistic regression can be achieved by starting with a standard implementation and applying FedAvg, which aggregates the local models after a specified number of iterations performed by the local logistic regressions.
5 Legal aspects of federated learning
The legal frameworks used within FL consortia are rarely discussed in the literature. Ballhausen et al. (2024) describes both the technical and legal aspects of a European pilot study implementing a federated statistical analysis by secure multiparty computation. They set up agreements between parties similar to those between the participants of a multi-centre clinical trial because using SMPC and only exchanging model gradients was legally considered data pseudonymisation (rather than anonymisation). FL was determined to require the same level of data protection as regular data sharing, which is also the most conservative course of action suggested in Truong et al. (2021); Lieftink et al. (2024). All clients jointly controlled the consortium and were responsible for determining the purpose and means of processing, including obtaining approval from the respective Ethics Committees. Sun et al. (2022a) similarly describes the server in their consortium as a trusted, secure environment supported by a legal joint controller agreement between the data owners.
Following Ballhausen et al. (2024), establishing an FL consortium could be expected to require all participating and involved parties to execute agreements that regulate their interactions. Doing so will establish the level of trust between parties and their responsibilities towards each other, third parties, and patients. Risk aversion suggests that it should contain a data-sharing agreement allowing for information sharing described in Figure 1. No party has access to other parties’ data. Still, it is theoretically possible that, in some cases, the model updates shared during FL could be deanonymised by malicious internal or external attackers (Truong et al., 2021). Parties may then be reluctant to treat that information as non-personally identifiable without formal mathematical proof of anonymisation and prefer to establish data protection responsibilities with a data-sharing agreement. In the EU, GDPR states that we should consider “all the means reasonably likely to be used” to “determine whether a natural person is identifiable” (European Union, 2016). Securing infrastructure in depth using best practices from information technology, defensive software engineering, and data by secure computing and encryption can make malicious attacks impractical with current technologies (security by design and by default; Volini, 2020; Martin and Kung, 2018). In addition, when FL involves models other than deep neural networks, if the contributions of individual parties are well balanced across the consortium and include a sufficiently large number of individuals, the information exchanged may very well be the same summary information routinely published as supplementary material to academic journal publications (Jégou et al., 2024). A recent systematic literature review of privacy attacks in FL has also highlighted that many of them are only feasible under unrealistic assumptions (Wainakh et al., 2022). Therefore, the reduced amount of information shared during FL and using secure computing should be considered to provide increased protection against data leaks and misuse. Advertising it as a key feature of the FL consortium will make partners and patients more comfortable with contributing to federated studies (see, for instance, Ballhausen et al., 2024). Lieftink et al. (2024), which investigated how FL aligns with GDPR in public health, also acknowledges that FL mitigates many privacy risks by enforcing purpose limitation, data use and information exchange minimisation, integrity and confidentiality (at a cost, as discussed in Section 2.4).
Furthermore, consortium parties must agree on how to assign intellectual property (IP) rights. Bioinformatics research often has practical applications in industry, which may involve patenting the results and apportioning any financial gains arising from their use. Parties in the consortium jointly control it and should share any gains from it (Minssen and Pierce, 2018). FL consortia are no different in this respect. From a technical standpoint, watermarking techniques for tracking data provenance and plagiarism have been adapted to FL Tekgul et al. (2021) to identify data and model theft.
In addition, the agreement establishing the FL consortium must set out its relationships with third parties and their legal duties. Third parties that have access to the infrastructure may be required to sign a data processing agreement to guarantee the safety and privacy of data. In many countries, patients also have the right to withdraw their consent to use their data, which in turn may require implementing procedures to remove individual data points from future federated analyses.
To summarise, FL provides increased protection against data and model leaks and should reduce the parties’ and patients’ perceived risk in contributing to the consortium. However, out of an abundance of caution, establishing a consortium-wide data-sharing agreement may allocate and reduce party responsibilities in the event of privacy breaches. The use of FL has a limited impact on other legal aspects of collaborative analysis, such as IP handling and requirements for third parties, because it is a technical solution that does not change the fundamental legal rights and responsibilities of the parties involved in the consortium.
6 Conclusions
Several independent research efforts have shown federated learning to be an effective tool to improve clinical discovery while minimising data sharing (Roth et al., 2020; Pati et al., 2022; Heyndrickx et al., 2023). FL allows access to larger and more varied data pools, leading to a faster and more robust exploration and translation of results. At the same time, it provides enhanced data privacy and can seamlessly incorporate advanced encrypted and secure computation techniques. Despite the increased computational requirements and reduced ability to explore and troubleshoot issues with the data (Lieftink et al., 2024), the benefits of FL may outweigh these additional costs.
Therefore, federated learning can potentially reduce the risks associated with EU and US regulations if implemented in a manner secure by design and by default. Its use may also make patients and institutions more confident in participating in clinical studies by reducing privacy and data misuse risks. However, the reliable use of federated learning and its effective translation into clinical practice requires a concerted effort by machine learning and clinical and information technology specialists. All their skills are required to accurately evaluate those risks and expand practical applications of federated learning in bioinformatics beyond the early-stage applications reviewed in this paper.
Funding
This work was supported by the European Union Horizon 2020 programme [101136962]; UK Research and Innovation (UKRI) under the UK Government’s Horizon Europe funding guarantee [10098097, 10104323] and the Swiss State Secretariat for Education, Research and Innovation (SERI).